OpenStudy (anonymous):

15(2 - x) = 13(3 - x)

OpenStudy (anonymous):

Multiply both sides out then put the x's on one side and the constants on the other side.

OpenStudy (anonymous):

Multiply them by what..? like 15x13? or?

OpenStudy (anonymous):

So whenever you see the parentheses it means that you have to mutiply the number that's outside of it by the numbers inside. I uploaded the solution.

OpenStudy (anonymous):

Alright, I get that. But the answer wasn't correct :/. I'm not sure what it's looking for.

OpenStudy (anonymous):

really? what answer do you have?

OpenStudy (anonymous):

-9/2 ?

OpenStudy (anonymous):

It shows that it's a fraction though.

OpenStudy (anonymous):

yeah -9/2 is my answer

OpenStudy (anonymous):

Alright, thanks for trying. I guess I'll keep looking for a solution/answer. :)

OpenStudy (anonymous):

I'm pretty sure my answer is correct lol..Do you have an answer to the problem there?

OpenStudy (anonymous):

I'll check what it says. Because everybody says it's -9/2.

OpenStudy (anonymous):

Ok good idea

OpenStudy (anonymous):

It says it's -4 1/2.

OpenStudy (anonymous):

:) that's the same as -9/2. Do you know how to simplify mixed numbers?

OpenStudy (anonymous):

Ohh, it wanted me to simplify. And i'm not entirely sure I do. Help? :)

OpenStudy (anonymous):

Sure. You say the whole number times the number at the bottom of the fraction (4*2) then add the number on top so in this case add 1. (8+1=9) Then you say the number you now have (9) divided by the number at the bottom of the fraction before (9)/(2) and don't forget the (-) sign you had before

OpenStudy (anonymous):

Thank you so much for your time and helpfulness!! :)

OpenStudy (anonymous):

No prob. i hope you understand it better now! :)

OpenStudy (anonymous):

:D You're awesome!

OpenStudy (anonymous):

Haha thanks :D

OpenStudy (anonymous):

I have a couple more questions. A bit more complicated! aha.

OpenStudy (anonymous):

Ok I'll take a look at them

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

A) 12(x -6) = 2(x - 9) B) 4(7 - 3x) = 7(4 - 2x) C) 7(x - 2) + 3(x + 2) = 5(x - 6) D) 4(2x -3) + 8(x - 4) = 2(2x + 6) E) 4(7x - 1) = 3(2x + 1) - 4(x + 6) F) 12(2x - 8) = 4(5 - 2x) + 3(7 - 3x) G) 3(x - 2) + 4(2x - 6) = 6(x - 4) + 8(2x + 1) Any help you could provide to giving or helping me find these answers would be lovely! :)

OpenStudy (anonymous):

What the other ppl said when you asked the question is totally right you know. i checked their work. :)

OpenStudy (anonymous):

Alright, thanks!

OpenStudy (anonymous):

no prob